# Exploring the Adversarial Vulnerabilities of Vision-Language-Action Models in Robotics

### v2

@@ 3.3 Untargeted Position-aware Attack
4:     $(x,y)\sim\mathcal{X}$
5:     for $i=1,2,\cdots,k$ do
6:        $shx,shy\sim\mathcal{U}(-\phi,\phi)$, $\theta\sim\mathcal{U}(-\psi,\psi)$
@@
[…]
10:        Clip $\delta$ into normal pixel range.
11:     end for
12:  end for
@@
3.4 Targeted Manipulation Attack
In addition to the aforementioned untargeted attacks, we also explore the adversarial vulnerabilities of VLA models with a targeted strategy. This attack aims to mislead the models into predicting specific trajectories, resulting in precise alterations. These targeted alterations can potentially drive malicious behaviors or induce task failure. The objective of the targeted manipulation attack is:
$\small\mathcal{L_{\text{TMA}}}=\mathbb{E}_{(x,y)\sim\mathcal{X}}[\mathcal{H}(\mathcal{F}(x+\delta),y^{I}_{T})],$ (7)
where $y^{I}_{T}=\{y^{i}=t|i\in[1,\dots,7],t\in\mathcal{R}\}$ represents the adversarial target; and $\mathcal{H}(\cdot,\cdot)$ denotes the Cross-Entropy loss, which measures the error between the predicted state under adversarial perturbation and the desired target state. By steering the robot towards an adversarial target across successive time steps, our approach manipulates the trajectory and undermines task performance, ultimately leading to a substantial disruption of intended operations. Furthermore, manipulating any individual DoF or a combination thereof can also severely compromise task success, given that the errors in one dimension tend to propagate and magnify over time during execution (see Fig. 3).
@@ 3.5 Implementation Details
The implementation of our pipeline is detailed in Algorithm 1. We incorporate two key modifications to enhance the training stability of the generated patch.
Inner-loops. Following previous work [51], we add $k$ inner-loops optimize steps during each iteration in line 5, aiming to reduce data variance and ensure more consistent updates.
Geometric Transformations. To improve the robustness of our attack under real-world scenarios, we employ random geometric transformations $\mathcal{T}(\cdot)$ in line 1 with affine and rotation transformations [3] as:
$\small\mathcal{T}(\cdot)=\begin{bmatrix}1&sh_{y}\\ sh_{x}&1\end{bmatrix}\cdot\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix},$ (8)
@@
4 Experiments
In this section, we present the experimental design outlined in §4.1 and report the quantitative and qualitative results of UADA, UPA, and TMA in both simulation and physical-world scenarios (§4.2). We then analyze diagnostic experiments (§4.3) to evaluate the influence of key components. Additionally, we assess the robustness of our method against various defense mechanisms in §4.4 and conclude with a comprehensive systemic discussion in §4.5.
@@ 4.1 Experiment Setup
Simulation Tasks. We conduct simulation attacks on BridgeData V2 [72] and LIBERO [46]. BridgeData V2 [72] is a real-world dataset comprising $24$ diverse environments and $13$ distinct skills, such as grasping, placing, and object rearrangement, with a total of $60,096$ trajectories. LIBERO [46] is a simulation dataset designed to evaluate models across four distinct task types (Spatial, Object, Goal, Long), each containing $10$ tasks with $50$ human-teleoperated demonstrations. Notably, LIBERO-Long combines diverse objects, layouts, and long-horizon tasks, making them challenging for complex, multi-step planning, thus we choose LIBERO-Long to conduct UADA, UPA and TMA.
@@
Robot Setups. For the physical-world tasks, we adopt a robotic system consisting of a Universal Robots UR10e equipped with a Robotiq Hand-E Gripper to provide a 7DoF motion. The sensing system includes one RGB webcam in the fixed position with a shoulder view.
@@
Victim VLA Models. In our study, we select publicly available and current most influential VLAs as victim models for evaluation. Specifically, we focus on four variants of the OpenVLA model, each trained independently on different tasks within the LIBERO dataset (i.e., Spatial, Object, Goal, Long). To evaluate the effectiveness of our methods, we craft adversarial patches using two distinct generating setups. The first involves a model trained in a simulated environment using the LIBERO dataset [46] with the openvla-7B-libero-long variant [35]. The second uses a model trained on physical world data from the BridgeData v2 dataset [72] with the openvla-7B model [35]. This approach allows us to assess our methods on both simulated and real-world data. Subsequently, we evaluate the performance of generated adversarial patches on victim models (i.e., OpenVLA LIBERO variants) trained on different tasks to rigorously prove the robustness and effectiveness of our method. These models are trained with distinct data sources and task objectives. Additional results on transferability (e.g., LLaRA [44]) are included in §S1.
Attacking Details. Aligned with the former objective [69] aiming to mislead classifiers and combine with our framework, we design and conduct an untargeted attack. Following [35], we apply scaling and normalization for data preprocessing. The initial learning rate is set to $2e^{-3}$, with the AdamW optimizer [49] with a cosine annealing scheduler, including $20$ warm-up iterations. Training is conducted over $2e^{3}$ iterations with a batch size of $6$ and $50$ inner-loop steps (see Algorithm 1). The transformation parameters $a,b$ in Eq. 8 are set to $0.2$ and $30$ to simulate real-world perspective change. For all three objectives, considering the trade-off between stealth and performance observed in §S2, we set the patch size to $5\%$. For the hyperparameters $\alpha,\beta$ in Eq. 6, We conduct grid search and tune $\alpha$ from $[0.2,0.4,0.6,0.8]$ and $\beta$ from $[0.8,0.6,0.4,0.2]$. We select $\alpha=0.8$ and $\beta=0.2$ based on the validation set performance. Ablation studies $w.r.t.$ $\alpha$ and $\beta$ are presented in §S6.
@@
Table 1: Untargeted results (Simulation). We report Failure rates (%) and NAD (%) in LIBERO simulation. denotes the in-domain victim model and dataset aligned with the patch generation model and dataset, denotes a transfer attack task involving a distinct victim model and dataset (see §4.1). The naive FR of each victim model in the corresponding task is: Spatial (15.3%), Object (11.6%), Goal (20.8%) and Long (46.3%). The FR ($\uparrow$) is highlighted in best and second best. Same for Tab. 2.
Victim models Objective Attack Actions NAD (%) Spatial Object Goal Long Untarget DoF1 14.1 100.0 100.0 100.0 100.0     DoF1-3 11.4 100.0 100.0 100.0 100.0 UADA DoF1 18.1 100.0 100.0 100.0 100.0     DoF1-3 14.3 100.0 100.0 100.0 100.0 UPA     DoF1-3 14.5 100.0 83.4 93.4 96.7
Table 2: Untargeted results (Physical). Transfer results to 4 victim models with adversarial patches generated from physical data and corresponding model.
@@
Victim model Objective Attack Actions NAD (%) Spatial Object Goal Long Untarget DoF1 26.7 85.4 56.4 77.3 65.8     DoF1-3 23.8 78.1 46.7 52.4 56.7 UADA DoF1 31.9 73.4 33.5 41.7 62.3     DoF1-3 26.4 85.9 49.5 32.1 60.5 UPA     DoF1-3 26.9 86.4 42.7 53.3 61.2
Evaluation Details. To assess the effectiveness and robustness of our methods, we conduct evaluations on the LIBERO dataset [46]. Each task suite is evaluated with $10$ trials, with fixed patch location (see §S5) This setup is for avoiding the patch obstructing task-related objects.
@@
Building on the concept of robotic tasks Success Rate (SR) introduced in LIBERO [46], we adopt Failure Rate (FR), defined as $1-\text{SR}$, as the primary evaluation metric for this study. To further quantify action discrepancy, we employ NAD (see Eq. 3) for untargeted attacks, as it considers relative distances to quantify deviation from the optimal action. A higher NAD value indicates a greater discrepancy between the ground truth and the predicted action. For targeted attacks, we utilize the L1 distance, which measures the absolute deviation from the specified target action.
Reproducibility. Our methods are implemented in Pytorch [58]. Experiments are conducted on a single NVIDIA A100-40GB GPU. Our code and attack video is available at https://anonymous.4open.science/r/roboticAttack-56FF.
@@
Table 3: Targeted Manipulation Attack results. Average failure rate (%) are reported in 4 LIBERO simulation tasks. The performance of AFR ($\uparrow$) is marked in best and second best. The initial AFR across the four tasks is 23.5%.
Generation Setup Metric Target Actions Simulation / DoF1 DoF2 DoF3 DoF4 L1 0.022 0.029 0.040 0.029 AFR 97.8 97.5 96.8 73.4 / … DoF5 DoF6 DoF7 DoF1-3 L1 0.024 0.023 0.046 0.031 AFR 96.8 88.3 84.1 95.5 Physical / DoF1 DoF2 DoF3 DoF4 L1 0.013 0.022 0.041 0.031 AFR 89.1 85.3 57.2 48.8 / … DoF5 DoF6 DoF7 DoF1-3 L1 0.032 0.018 0.049 0.027 AFR 54.3 46.2 57.8 86.8
@@ 4.2 Main Result
Quantitative Results. For UADA and UPA, our methods effectively amplify action discrepancies, leading to a notable transfer attack ability in increasing failure rates (see Tab. 1). Specifically, for the Simulation setup, UADA and UPA achieve maximum NAD of 18.1% and 14.5%, significantly outperforming untargeted scenarios with increment of 3.8% and 0.4%, respectively. Both untargeted and UADA successfully disrupt the robot execution, with a failure rate to 100%. UPA also results in a significant increment compared to the failure rate of the victim methods in benign scenarios, which is 100% $v.s.$ 15.3%, 83.4% $v.s.$ 11.6%, 93.4% $v.s.$ 20.8% and 96.7% $v.s.$ 46.3%, respectively. Regarding the Physical setup, both UADA and UPA generate actions with large action discrepancies, which are 31.9% and 26.9% with an increase of 5.2% and 0.2% compared to the best 26.7% in untargeted scenarios, respectively. UPA demonstrates an improvement of 86.4% $v.s.$ 85.9% over the spatial task compared to untargeted.
@@
For TMA task, we evaluate the effectiveness of our method by manipulating various DoF to 0. Our method demonstrates significant effectiveness in manipulating robotic trajectory and increasing failure rates (see Tab. 3). Specifically, when applied to the different generation settings, our approach yields notable increments across various tasks, including an max average failure rate 97.8% $v.s.$ 23.5% of benign performance in Simulation, and 89.1% in Physical, respectively. To further prove the generalizability, we target attack on several non-zero values in §S2.
@@
Qualitative Results. We qualitatively analyze the trajectories of robot movement of the proposed three different objectives in Fig. 3. For UADA, we compare the trajectory deviations in the same trail with patch generated from UADA and untarget attack. As seen, the untarget attack induces small deviations in the trajectory UADA, on the other hand, produces significantly larger trajectory deviations (also supported by NAD metric in Tab. 2), which is attributed to UADA’s capability of incorporating action discrepancies. In our observation, UADA significantly amplifies the effect on overall task execution, thereby increasing the potential for robotic hazards. For UPA, we observe chaotic and irregular behaviors, including instances where the end-effector moves out of the camera’s field of view. We attribute this phenomenon to the efficacy of the adversarial patch in perturbing the model’s spatial perception, inducing a consistent deviation from the intended direction of movement, ultimately resulting in failure in a long run.
For TMA, we observe a notable reduction in the range of motion along the x-axis corresponding to the targeted attack axis. This suggests that our proposed targeted attack can effectively constrained robot movements.
@@
Physical-world Performance. In addition to the digital simulation results, we also conduct a comprehensive evaluation of the performance of our generated adversarial patches in physical-world testing scenarios. The evaluation encompassed 100 trials across three distinct tasks: object grasping, placement, and manipulation (see §S3 for more details). As shown in Fig. 5, the adversarial patch generated with UADA demonstrated the ability to manipulate the robot effectively, achieving an attack success rate exceeding 43%. Although this success rate is lower than the corresponding digital-world performance (i.e., 85.9%), it highlights the effectiveness of our patches in physical-world applications as well without the need for further adaptations. Crucially, we observed that the induced erratic movements (similar to simulation scenario) of the robotic manipulators during successful attacks pose significant risks to human safety and surrounding environments. This finding emphasizes the severe threat posed by VLA models in real-world applications, consistent with similar observations in digital settings.

### v3

@@ 4 Experiments
4:     for $i=1,2,\cdots,k$ do
5:        $sh_{x},sh_{y}\sim\mathcal{U}(-\phi,\phi)$, $\theta\sim\mathcal{U}(-\psi,\psi)$
6:        $y_{\text{pred}}\leftarrow\mathcal{F}\big{[}\mathcal{T}((x+\delta),(sh_{x},sh_{y},\theta))\big{]}$
@@
8:     end for
9:  end for
@@ 4.1 Implementation Details
The implementation of our adversarial patch attack pipeline is detailed in Algorithm 1. We incorporate two key modifications to enhance the training stability of the generated patch.
Inner-loops. Following previous work [49], we add $k$ inner-loops optimize steps during each iteration in line 4, aiming to reduce data variance and ensure more consistent updates.
Geometric Transformations. To improve the robustness of our attack under real-world scenarios, we employ random geometric transformations $\mathcal{T}\big{[}(\cdot),(sh_{x},sh_{y},\theta)\big{]}$ in line 1 with affine and rotation transformations [3] as:
$\small\mathcal{T}\big{[}(\cdot),(sh_{x},sh_{y},\theta)\big{]}=\begin{bmatrix}1&sh_{y}\\ sh_{x}&1\end{bmatrix}\cdot\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix},$ (9)
@@
Untargeted Manipulation Attack (UMA) Baseline. As a pioneering work, we could not find a directly comparable baseline for our attacks. Therefore, we adapt prior work in adversarial learning as one of our baseline methods [66]. This attack is designed to mislead classifiers; we integrate it into our framework, leading to the following objective:
$\small\mathcal{L_{\text{UMA}}}=-\mathbb{E}_{(x,y)\sim\mathcal{X}}[CE(\mathcal{F}(x+\delta)^{I},y^{I}_{gt})].$ (10)
By minimizing the negative CE loss, this attack pushes the predicted action away from the ground truth action.
Random Noise Baseline. We generate random noise patches as an additional baseline, representing an unstructured intervention without any learned adversarial intent. Specifically, the noise is sampled from a Gaussian distribution $\mathcal{N}(0,1)$ and evaluated with the same setting as proposed attacks.
@@ 4.2 Experiment Setup
Datasets. We conduct attacks on BridgeData V2 [69] and LIBERO [44] with corresponding VLAs. BridgeData V2 [69] is a real-world dataset comprising $24$ diverse environments and $13$ distinct skills, such as grasping, placing, and object rearrangement, with a total of $60,096$ trajectories. LIBERO [44] is a simulation dataset designed to evaluate models across four distinct task types (Spatial, Object, Goal, Long). Notably, LIBERO-Long combines diverse objects, layouts, and long-horizon tasks, making them challenging for complex, multi-step planning, thus we choose LIBERO-Long to conduct UADA, UPA and TMA.
@@
Robot Setups. For the real-world tasks, we adopt a robotic system consisting of a Universal Robots UR10e equipped with a Robotiq Hand-E Gripper to provide a 7DoF motion. The sensing system includes one RGB webcam in the fixed position with a shoulder view.
@@
Victim VLAs. In our study, we select publicly available and current most influential VLAs as victim models for evaluation. Specifically, we focus on four variants of the OpenVLA model, each trained independently on different task suites within the LIBERO dataset [44] (i.e., Spatial, Object, Goal, Long). To evaluate the effectiveness of our methods, we craft adversarial patches using two distinct generating setups: Simulation Setting involves a model trained in a simulated environment using the LIBERO Long suite with the openvla-7B-libero-long variant [34]. Physical Setting uses a model trained on physical world data from the BridgeData v2 [69] with the openvla-7B model [34]. This approach allows us to assess our methods on both simulated and real-world data. Subsequently, we evaluate the performance of generated adversarial patches on victim models (i.e., OpenVLA LIBERO variants) trained on different tasks suites to rigorously prove the robustness and effectiveness of our method. These models are trained with distinct data sources and task objectives. Additional results on transferability (e.g., LLaRA [42]) are included in §S2.
@@
Table 1: Untargeted results (Simulation). We report Failure Rates (FR) and NAD in LIBERO simulation. denotes the in-domain victim model and dataset aligned with the patch generation model and dataset, denotes a transfer attack task involving a distinct victim model and dataset (see §4.2). The FR ($\uparrow$) is highlighted in best and second best for each task suite. Same for Tab. 2.
@@
Victim Model Objective Action(s) NAD(%) Spatial Object Goal Long Avg Benign - - 15.3$\pm$10.211.6$\pm$10.020.8$\pm$12.046.3$\pm$18.623.5% Random Noise 28.8$\pm$24.214.8$\pm$7.921.0$\pm$15.548.4$\pm$14.828.3% UMA DoF1 14.1 100$\pm$0.099.0$\pm$3.0100$\pm$0.0100$\pm$0.099.8%     DoF1∼3 11.4 100$\pm$0.098.8$\pm$2.699.0$\pm$2.4100$\pm$0.099.5% UADA DoF1 21.0 100$\pm$0.099.2$\pm$2.4100$\pm$0.0100$\pm$0.099.8%     DoF1∼3 17.9 100$\pm$0.0100$\pm$0.0100$\pm$0.0100$\pm$0.0100% UPA     DoF1∼3 14.5 96.2$\pm$11.477.8$\pm$12.588.0$\pm$10.496.8$\pm$3.089.7%
Evaluation Details. To assess the effectiveness and robustness of our methods, we conduct experiments on the LIBERO dataset [44]. Each suite consists of 10 tasks, with each task executed for 50 trials, resulting in a total of 500 rollouts, following Kim et al. [34]. For reproducibility, we carefully select patch paste locations for each suite, ensuring they do not obscure objects or robots in the test environment. Further details are provided in §S3.
@@
Evaluation Metric. Regarding the task execution evaluation, we take the maximum steps of each task suite in the LIBERO training dataset as the timeout failure condition to reduce computational overhead. Furthermore, building on the concept of Success Rate (SR) introduced in LIBERO [44], we adopt Failure Rate (FR), defined as $\small 1-\text{SR}$, as the primary evaluation metric. To further quantify action discrepancy, we employ NAD (see Eq. 8) for untargeted attacks on corresponding DoF(s), as it considers relative distances to quantify deviation from the optimal action. A higher NAD value indicates a greater discrepancy between the ground truth and the predicted action. Additionally, we report the standard deviation of FR across tasks within each task suite.
Table 2: Untargeted results (Physical). Transfer results to 4 victim models with adversarial patches generated from physical settings. We report Failure Rates (FR) and NAD. FR ($\uparrow$) and NAD ($\uparrow$) are highlighted in best and second best for each task suite.
Victim Model Objective Action(s) NAD(%) Spatial Object Goal Long Avg Benign - - 15.3$\pm$10.211.6$\pm$10.020.8$\pm$12.046.3$\pm$18.623.5% Random Noise 28.8$\pm$24.214.8$\pm$7.921.0$\pm$15.548.4$\pm$14.828.3% UMA DoF1 26.7 89.6$\pm$15.360.8$\pm$16.864.8$\pm$26.880.0$\pm$20.873.8%     DoF1∼3 23.8 96.6$\pm$6.256.4$\pm$20.480.0$\pm$20.582.0$\pm$17.978.8% UADA DoF1 32.6 99.2$\pm$1.898.8$\pm$2.492.6$\pm$15.296.6$\pm$4.896.8%     DoF1∼3 27.7 100$\pm$0.0100$\pm$0.0100$\pm$0.0100$\pm$0.0100% UPA     DoF1∼3 26.9 95.6$\pm$9.5 % 57.0$\pm$23.657.2$\pm$24.472.8$\pm$19.770.7%
@@
Table 3: Targeted Manipulation Attack results. Failure Rate (FR, $\uparrow$) and its standard deviation across tasks within LIBERO [44] suite are reported. The performance of task suits and DoFs is averaged separately, with the best and second-best values separately.
Simulation (Attacked on OpenVLA-LIBERO-10) Physical (Attacked on OpenVLA-7b) Targeted DoF(s) Metric Spatial Object Goal Long Avg Spatial Object Goal Long Avg Benign FR 15.3$\pm$10.211.6$\pm$10.020.8$\pm$12.046.3$\pm$18.623.5% 15.3$\pm$10.211.6$\pm$10.020.8$\pm$12.046.3$\pm$18.623.5% Random Noise 28.8$\pm$24.214.8$\pm$7.921.0$\pm$15.548.4$\pm$14.828.3% 28.8$\pm$24.214.8$\pm$7.921.0$\pm$15.548.4$\pm$14.828.3% DoF1 FR 100$\pm$0.097.8$\pm$3.0100$\pm$0.0100$\pm$0.099.5% 90.6$\pm$19.936.6$\pm$18.161.0$\pm$21.786.4$\pm$12.768.7% DoF2 100$\pm$0.099.2$\pm$1.3100$\pm$0.0100$\pm$0.099.8% 99.6$\pm$0.869.2$\pm$11.176.2$\pm$26.191.8$\pm$9.984.2% DoF3 100$\pm$0.097.2$\pm$6.0100$\pm$0.0100$\pm$0.099.3% 72.0$\pm$14.830.2$\pm$17.439.2$\pm$20.962.4$\pm$17.551.0% DoF4 94.0$\pm$11.037.4$\pm$23.251.2$\pm$26.587.6$\pm$14.967.6% 93.6$\pm$9.731.8$\pm$20.345.0$\pm$18.363.8$\pm$17.058.6% DoF5 100$\pm$0.092.6$\pm$8.4100$\pm$0.0100$\pm$0.098.2% 97.8$\pm$4.936.0$\pm$22.562.0$\pm$24.879.4$\pm$21.368.8% DoF6 100$\pm$0.080.0$\pm$9.592.6$\pm$13.4100$\pm$0.093.2% 79.0$\pm$14.222.2$\pm$14.441.0$\pm$21.562.6$\pm$15.751.2% DoF7 99.2$\pm$1.386.6$\pm$8.367.2$\pm$22.596.4$\pm$6.487.4% 91.8$\pm$15.398.8$\pm$1.370.6$\pm$23.493.2$\pm$10.288.6%  DoF1∼3 100$\pm$0.0100$\pm$0.0100$\pm$0.0100$\pm$0.0100% 96.4$\pm$8.083.6$\pm$16.874.4$\pm$26.391.8$\pm$11.986.6% Task Avg 99.2% 86.4% 88.9% 98.0% - 90.1% 51.1% 58.7% 78.9% -
@@ 4.3 Main Result
Quantitative Results. For UADA and UPA, our methods effectively amplify action discrepancies, leading to a notable transfer attack ability in increasing failure rates (see Tab. 1). Specifically, while attacking DoF1 and DoF1∼3 in the Simulation setup, UADA and UPA achieve NAD of 21.0% and 14.5%, significantly outperforming UMA scenarios with increments of 6.9% and 3.1%, respectively. Both UADA and UPA effectively disrupt robot execution, yielding maximum average failure rates of 100% and 89.7%, respectively. Regarding the Physical setup (see Tab. 2), our attack methods demonstrate strong transferability, as adversarial patches generated in the physical setting significantly impact the performance of LIBERO-X test. Specifically, both UADA and UPA generate malicious actions with large action discrepancies, which are 32.6% (DoF1) and 26.9% (DoF1∼3) with an increase of 5.9% and 3.1% compared to 26.7% (DoF1) and 23.8% (DoF1∼3) in UMA scenarios, respectively. Compare the two settings (Tab. 1 $v.s.$ Tab. 2), we observe a general variance in NAD (e.g. 32.6% $v.s.$ 21.0% for UADA). This discrepancy can be attributed to the fundamental differences between the simulation and real-world datasets used for conducting attacks (Bridge V2 [69] $v.s.$ LIBERO [44]). The increased variability in real-world data, including environmental complexity, object diversity, and task difficulty, allows the robot more opportunities to generate larger action discrepancies within the validation dataset.
@@
For TMA task, we evaluate the effectiveness of our method by manipulating all DoF(s) to 0 (i.e., $y^{i}_{T}=0$ in Eq. 6). Our method demonstrates significant effectiveness in manipulating robotic trajectory and increasing FR (see Tab. 3). Specifically, when applied to the different generation settings, our approach yields notable increments across various tasks, including a max average failure rate 100% $v.s.$ 23.5% of benign performance in Simulation, and 88.6% in Physical, respectively. Moreover, during attack DoF4, we observe low FR and high task deviation. This failure can be attributed to the fact that DoF4 controls the orientation along the x-axis, which can be redundant DoF in tasks. As a result, attacks targeting redundant DoF(s) [38] are less likely to disrupt execution effectively. This observation underscores the importance of task-specific considerations when designing adversarial attacks on robotic systems. To further prove the generalizability, we target on several non-zero values in §S4.
@@
Qualitative Results. We qualitatively analyze robot movement trajectories under the three proposed attacks in Fig. 3. For UADA, we compare the trajectory deviations in the same trail with patches generated from UADA and UMA attacks. As seen, the UMA induces small deviations in the trajectory. UADA, on the other hand, produces significantly larger trajectory deviations (also supported by NAD metric in Tab. 2), which is attributed to UADA’s capability of incorporating action discrepancies. In our observation, UADA significantly amplifies the effect on overall task execution, thereby increasing the potential for robotic hazards. For UPA, we observe chaotic and irregular behaviors, including instances where the end-effector moves out of the camera’s field of view. We attribute this phenomenon to the efficacy of the adversarial patch in perturbing the model’s spatial perception, inducing a consistent deviation from the intended direction of movement, ultimately resulting in failure in a long run. For TMA, we observe a notable reduction in the range of motion along the x-axis corresponding to the targeted attack axis. This suggests that our proposed targeted attack can effectively constrained robot movements.
@@
Real-world Performance. In addition to the digital simulation results, we also conduct a comprehensive evaluation of the performance of our generated adversarial patches in real-world scenarios. The evaluation encompassed 100 trials across three distinct tasks: object grasping, placement, and manipulation (see §S5 for more details). As shown in Fig. 4, the adversarial patch generated with UADA demonstrated the ability to manipulate the robot effectively, achieving an attack success rate exceeding 43%. Although this success rate is lower than the corresponding digital-world performance (i.e., 100%), it highlights the effectiveness of our patches in physical-world applications as well without the need for further adaptations. Crucially, we observed that the induced erratic movements (similar to simulation scenario) of the robotic manipulators during successful attacks pose significant risks to human safety and surrounding environments. This finding emphasizes the severe threat posed by VLA models in real-world applications, consistent with similar observations in digital settings.